# Non-reciprocal neutral ferroelectric domain walls in BiFeO3

### v1

@@ Abstract
This paper analyzes a peculiar phenomenon of non-reciprocal domain wall pairs and illustrate the implications in ab-initio-based atomistic computational experiments with ($\bar{1}\bar{1}2$)-oriented planar R180 domain walls within the canonical multiferroic ferroelectric crystal of BiFeO3. Results show that parallel walls on the opposite sides of a given domain within a simply twinned lamellar domain structure can have considerably different polarization and oxygen octahedra tilt profiles, different thickness and energy densities. The spontaneous formation of zigzag walls and triangular domains suggests that these domain walls are actually the lowest energy planar R180 walls in pure insulating BiFeO3 crystals and films.
@@
Domain structure is a natural manifestation of symmetry breaking phase transitions in crystals. One of the simplest but still very frequent domain structure is the simple lamellar twinning, consisting in alternation of only two distinct domains states, $A$ and $B$, separated by parallel planar domain walls. Often, the two subsequent walls $A|B$ and $B|A$ are equivalent by symmetry reasons. However, it is also possible that the symmetry relationship between A||||B and B||||A is absent. Then, $A|B$ and $B|A$ are physically non-equivalent walls. Such non-reciprocal walls were earlier also referred as irreversible onesJanovec .
@@
The non-reciprocity applies to all nominally charged ferroelectric domain walls. Such walls show a mismatch in the normal component of the spontaneous polarization in the adjacent domains. In lamellar case, the two subsequent charged walls can be denoted head-to-head (HH) and tail-to-tail (TT) wall, respectively. The HH and TT walls differ by the sign of the interfacial charge and thus the absence of symmetry relation is obvious. For example, it was reported that the electric conductivity of HH and TT 90 walls in barium titanate differs by several orders of magnitude.
@@
In this paper, we are addressing the question whether and when one can encounter the manifestations of a much less obvious non-reciprocity of nominally neutral ferroelectric walls. In particular, we explore the case of the experimentally reported but so far little investigated $(\bar{1}2\bar{1})$ oriented 180 walls in BiFeO3 and argue that these walls are indeed showing remarkable non-reciprocal effects.
BiFeO3 is a well-studied multiferroic material with coexisting ferroelectric, ferroelastic, antiferrodistortive and antiferromagnetic orders and promising applications in spintronics, nonvolatile memory devices, catalysis and sensors wang_2020 ; chaudron_2024 ; YangAbove ; Amdouni . At room temperature, it has rhombohedral $R3c$ symmetry with a spontaneous polarization between $0.6-1.0~{}C/m^{2}$ and with oxygen octahedra alternatively tilted by about $13.8^{\circ}~{}$ around the polarization axis  lebeugle_2007 ; Catalan09 . The conventional hexagonal unit cell is depicted in Fig. 1. The parent cubic perovskite $m\bar{3}m$ reference phase gives rise to eight equivalent polarization states. Corresponding ferroelectric domain walls are primarily distinguished as ${}^{\circ}~{}$, ${}^{\circ}~{}$ and ${}^{\circ}~{}$ ones according to the angles between the polarization in the adjacent domains. Details structure and properties of domain walls in BiFeO3 have been subject of many experimentalFarokhipoor ; Sando14 ; Bhatnagar and theoretical studiesDieguez2013 ; Ren2013 ; ourPRL . The benchmark systematic first-principles study of Ref. Dieguez2013, explored almost all types of mechanically and electrically compatible (neutral) domain walls. However, non-reciprocal walls were not addressed there. The present work fills this knowledge gap and employ an atomistic model to demonstrate and explain several remarkable non-reciprocal domain wall effects with specific implications for this important model ferroelectric material.
To obtain 180 domain walls we used an interatomic potential with parameters fitted from ab-initio calculations Graf2014 ; Graf2014b that has been also employed in Refs. Graf2015, ; Graf2016, ; ourPRL, . The potential utilizes the framework of the shell model, representing each atom in the system with a core and a shell, mimicking the atomic polarizability and with the anharmonic core-shell interaction is particularly well-suited in reproducing ferroelectric behavior  Tinte2004 . The low-temperature multidomain configuration was stabilized using a supercell made of 20$\times$20$\times$2 rhombohedral BiFeO3 cells in the hexagonal setting (i.e. 24000 atoms) subjected to periodic boundary conditions.
The structural relaxation was achieved using classical molecular dynamics simulations under constant stress and temperature conditions. A temperature of 1 K was chosen to facilitate effective relaxation while introducing a minimal amount of thermal noise. The simulation began with a 10 ps thermalization, followed by an additional 10 ps during which the average configuration was computed. The molecular dynamics calculations were performed using the DL_POLY software  Todorov2006 . Local dipole moments were evaluated for each 5-atom perovskite cell centered in the Bi atoms, based on the positions and charges of cores and shells. For a given cell, one Bi atom with its surrounding twelve O atoms (with a weight of 0.25) and eight coordinating Fe atoms (0.125 weight) were considered.
@@
The present study focuses on ${}^{\circ}~{}$ domain walls. Keeping the usual notation where the Miller indices $h,k,l$ refer to the parent cubic structure. We choose two antiparallel domain states with polarization along $[111]$ and $[\bar{1}\bar{1}\bar{1}]$. The mechanically and electrically compatible ${}^{\circ}~{}$ domain walls connecting these two states coincide with planes parallel to the $[111]$ direction, that we also identify with positive sense of $z^{\prime}$ axis (see Fig. 1a). There are two high-symmetry options for such a plane; $(1\bar{1}0)$-type planes, parallel to the mirror symmetry plane of the parent $m\bar{3}m$ reference phase, and ($\bar{1}\bar{1}2$)-type planes, passing through the intersecting 3-fold and 2-fold rotational axes of the parent $m\bar{3}m$ reference phase. Most of the studies have considered only the former type, but as we shall argue, the ($\bar{1}\bar{1}2$)-type walls are even more important for BiFeO3.
The desired ($\bar{1}\bar{1}2$) domain wall orientation was selected by imposing the suitable step-like distribution of the initial Bi ion off-centering displacements in our simulation box. In order profit from the computational advantages of the periodic boundary conditions, one has to introduce two such domain walls, as indicated in Fig. 1b.
After the relaxation of the structure, most of the volume is in the equilibrium single domain structure with the spontaneous polarization either parallel or antiparallel to the sense of the $z^{\prime}$ axis. For the detailed inspection of the domain walls, polarization of each unit cell was projected on the rotated coordinate system, where $s\parallel[\bar{1}\bar{1}2]$ is perpendicular to the wall, $z^{\prime}\parallel[111]$ is along the polarization, and $t\parallel{\bf b_{\rm h}}$ is perpendicular to the both $s$ and $z^{\prime}$.
@@
Resulting polarization profiles in the vicinity of domain wall DW1 and DW2 are shown in Fig. 2a,b. Analogous profiles of the staggered oxygen octahedra tilt rotations $R_{i}$ are shown in Fig. 2cd. As expected, in both DW1 and DW2 the $P_{\rm s}$ and $P_{\rm t}$ components interpolate between zero values, while the $P_{\rm z^{\prime}}$ component is switching between the nominal spontaneous polarization values. The absence of switching of the $R_{\rm i}$ components confirms that neither DW1 nor DW2 coincide with an anti-phase boundary of the underlying staggered tilt pattern and thus both walls can be considered as purely ferroelectric domain walls.
@@
While the profiles of $P_{\rm s}$, $P_{\rm z^{\prime}}$, $R_{\rm s}$, and $R_{\rm z^{\prime}}$ components are same at any line passing through a given wall, the values of $P_{\rm t}$ and $R_{\rm t}$ alternate between two mutually opposite values when comparing lines related by the fractional translation vector ${\bf c_{\rm h}}/2$. It means that the Bloch polarization component averaged over the two sites related by ${\bf c_{\rm h}}/2$ is zero, and at the same time, there is a staggered (antiferroelectric) Bloch component within the wall. In case of the tilt profile, the double signs of $R_{\rm t}$ implies that the Bloch component of the staggered oxygen octahedra rotation averaged over the two octahedra related by ${\bf c_{\rm h}}/2$ is zero, while both octahedra are slightly rotated in the same sense around the $t$-axis.
@@
However, there are also important differences between DW1 and DW2. First, the DW1 wall has thickness of about 0.7 nm while DW2 thickness is about 0.35 nm Second, the magnitude of the tilt modulations are much stronger in DW1 and moreover, the signs of the $R_{\rm s}$ are opposite in DW1 and DW2. A mode detailed analysis of the stuctural profiles within a continuous Ginzburg-Landau theory is a subject of a separate report.
@@
In order to verify that the resulting profiles shown in Fig. 2 are the lowest energy domain wall configurations, we have slightly varied the initial positions of the walls, size of the simulation box and we also annealed at higher temperatures. In overall, the domain characteristics described proved to be robust and representative ones.
Nevertheless, a very interesting phenomenon was encountered in the course of gradual temperature increase in steps of 20 K up to 300 K. This computer experiment, documented in Fig. 3, has been performed in the supercell made of $20\times 20\times 8$ rhombohedral hexagonal cells (i.e. 96000 atoms). The molecular dynamics trajectory for each temperature began with a 10 ps thermalization, followed by an additional 10 ps for computing temporal averages from snapshots recorded every 0.05 ps. The final configuration of the previous temperature was used as the starting point for the next temperature.
The time-averaged polarization map at $z^{\prime}=0$ was practically same till $T=120$ K (see Fig. 3a). However, when the temperature reached $T=140$ K, the DW2 started to display significant fluctuations. The original planar wall changed to an irregularly corrugated shape and the $P_{\rm t}$ components started to be disordered. The slow fluctuations are apparent even in the 10 ps averaged polarization maps (see Fig. 3b). Finally, at $T=240$ K, the habitus of DW2 abruptly transformed into a zigzag wall with ${}^{\circ}~{}$ folds. This form of DW2 then persists till $T=300$ K (see Fig. 3c) and it remains there even when they system is cooled back down to $T=1$ K.
@@
Finally, we also tried to stabilize small diameter columnar nanodomain of inverted polarization, still with neutral domain walls, but otherwise without any constraint from the periodicity of boundary conditions in the $x^{\prime}y^{\prime}$ plane. Obviously, small wall segments or the curved domain walls are more strongly interacting but they are also more free to change their position or orientation. For a $P_{\rm z^{\prime}}>0$ nanodomain, the only encountered stable configurations had a cross section of an equilateral triangle, oriented as shown in Fig. 4. We could see that one of the sides of this right triangular prism is always parallel the walls of Fig. 1b and that the wall profiles correspond to that of DW1 in Fig. 2ac. Moreover, we never seen $(1\bar{1}0$-type walls suggesting that DW1 is actually the lowest energy 180 domain wall.
@@
Summarizing, we have shown that the DW2 differs from DW1 the thickness, amplitudes of the secondary order parameters in the walls and also by the respective stability with respect to temperature and domain wall interactions. The ensemble of all these findings demonstrates that DW1 and DW2 are not related by any symmetry operation.
@@
Figure 4: Simulations of triangular nanodomains. (a)Polarization map in of the relaxed structure of a triangular nanodomain. (b) The excess energy due to the domain wall as a function of the lateral extension of the nanodomain and the linear fit to the data.
@@
For a more precise discussion, let us introduce a more instructive domain wall symbols $A|{\bf{n}}\rangle B]$. Here $A$ and $B$ stands for domain states while the oriented domain wall normal $\bf{n}$ points in the real space towards the domain that is ascribed to the state of the right-hand side of the symbol, as it is mnemonically indicated by the embedded ”ket” symbol.
Therefore, using the already introduced pseudocubic Miller indices for spontaneous polarization directions and oriented domain wall normals, our DW1 of Fig. 1b can be described as either $[111|1\bar{2}1\rangle\bar{1}\bar{1}\bar{1}]$ or $[\bar{1}\bar{1}\bar{1}|\bar{1}2\bar{1}\rangle 111]$, while DW2 can be only described as $[111|\bar{1}2\bar{1}\rangle\bar{1}\bar{1}\bar{1}]$ or $[\bar{1}\bar{1}\bar{1}|1\bar{2}1\rangle 111]$.
@@
Symmetry of planar domain walls was systematically developed by JanovecJanovec . The basic assumption is that specific symmetries are described by certain subgroups of the parent phase crystallographic group $G$. The proper symmetry of the $[A|{hkl}\rangle B]$ wall is described by a group $T_{AB}|{hkl}\rangle$, composed of those parent phase symmetry group operations, that results in an indistinguishable domain wall geometry. It is a crystallographic group with two-dimensional translational symmetry. The highest possible symmetry $T_{AB}(hkl)$ can be determined considering the orientation of the domain wall plane and its centering in the crystal structure. In the present case, the $2_{\rm t}/m_{\rm t}$ factor group of $T_{AB}(hkl)$ contains the identity, inversion $i$, mirror $m_{\rm t}$ perpendicular to $t$ axis and 2-fold rotation $2_{\rm t}$ parallel to the $t$ axis. Both planar DW1 and planar DW2 walls agree with this symmetry within the precision of our numerical calculations.
@@
The non-reciprocity is related to the symmetry group of domain wall pair $\bar{J}_{AB}|(hkl)$, which collects all parent phase symmetry operations that preserve or exchange the domain states and preserve the domain wall plane. In principle, this is supergroup of $T_{AB}(hkl)$ because it can also contain operations that interchange the sides of the wall but do not alter the domain state (pure side-reversal operations) or operations that reverse the domain states but do not interchange sides of the wall (pure state-reversal operations). Existence of pure side-reversal or pure state-reversal symmetry implies the reciprocity (reversibility) of the wall. In the case of our wall, no such symmetry operations exists and $\bar{J}_{AB}|(hkl)=T_{AB}(hkl)$.
@@
Finally, the group of the unordered domain pair, $J_{AB}$, involves all operations maintaining or interchanging the two domain states. In our case the $\bar{3}m$ factor group of $J_{AB}$ is a nontrivial supergroup of the factor group of $\bar{J}_{AB}|(hkl)$ what implies that there are the walls equivalent to DW1 that joins the same two domain states. It can be seen that these are the $[111|\bar{2}11\rangle\bar{1}\bar{1}\bar{1}]=[\bar{1}\bar{1}\bar{1}|\bar{1}2\bar{1}\rangle 111]$ and $[111|11\bar{2}\rangle\bar{1}\bar{1}\bar{1}]=[\bar{1}\bar{1}\bar{1}|\bar{1}\bar{1}2\rangle 111]$ walls corresponding to the segments of the the zigzag wall in Fig. 3b. All these three equivalent walls are realized in the three sides of the triangular prism in Fig. 4a.
The irreversible decomposition of DW2 wall documented in Fig. 3 and the systematic preference for DW1-type domain walls in stable prismatic nanodomains (Fig. 4a) suggests that the planar domain wall energy density $\sigma_{2}$ of DW2 is significantly higher then $\sigma_{1}$ of DW1.
@@
The broadly applied straightforward procedure determination of the domain wall energy density from computational experiments consists in calculation of energy of planar domains with desired crystallographic orientation in a periodic cell like that of Fig. 1, subtracting the corresponding energy of the single domain state, and dividing the difference by the total surface of the walls within the supercell. In case of non-reciprocal wall pair, however, this yields only the average energy density $(\sigma_{1}+\sigma_{2})/2$. In order to overcome this topological obstacle, we have thus used the prismatic nanodomain of Fig. 4a. As argued above, the all three boundaries of these nanodomains are equivalent to DW1. For elimination of the domain fold energy contributions, we have tuned the initial conditions in a way to obtain a whole sequence of similar relaxed nanodomain configurations but with the base edge $l$ ranging from 5 to 20 nm. The slope of the excess energy associated with the domain walls vs the base edge lenght $l$ shown in Fig. 4b yielded $\sigma_{1}\approx 40$ mJ/m2. In combination with the average energy density obtained from the standard configuration of Fig. 1, we could see that $\sigma_{2}\approx 130$ mJ/m2, which is indeed significantly larger than $\sigma_{1}$.
@@
In summary, we have explored the R180 domain walls within BiFeO3 and find out that so far little investigated ($\bar{1}\bar{1}2$)-oriented R180 walls are the lowest energy ones. We have given symmetry arguments why these walls are non-reciprocal and explained various profound consequences of this peculiar property.
@@
This work was supported by the Czech Science Foundation (project no. 19-28594X) and by the European Union’s Horizon 2020 research and innovation programme under grant agreement no. 964931 (TSAR). MAPG was supported by the European Union and the Czech Ministry of Education, Youth and Sports (Project: MSCA Fellowship CZ FZU I - CZ.02.01.01/00/22__\__010/0002906).

### v2

@@ Abstract
This paper analyzes a peculiar phenomenon of non-reciprocity of domain walls and illustrates its implications using ab-initio-based atomistic computational experiments with ($\bar{1}\bar{1}2$)-oriented planar 180-domain walls within the canonical multiferroic ferroelectric crystal of BiFeO3. Results show that walls on the opposite sides of a given domain lamellae within a twinned domain structure can have considerably different properties, such as different polarization and oxygen octahedra tilt profiles, different thicknesses and different wall energy densities. The spontaneous formation of 180 zigzag walls and triangular domains of inverted polarization suggests that one of the non-reciprocal ($\bar{1}\bar{1}2$)-oriented domain walls is actually the lowest energy 180 wall of the pure insulating BiFeO3.
@@
Domain structure is a natural manifestation of symmetry breaking phase transition in a crystal. A frequent domain pattern is the lamellar twinning, formed by alternation of two distinct domains states, $A$ and $B$, and separated by parallel planar domain walls. Often, the two subsequent walls $A|B$ and $B|A$ are equivalent by symmetry reasons. However, it is also possible that the symmetry relationship between A||||B and B||||A is absent. Then, $A|B$ and $B|A$ are physically non-equivalent walls. Such non-reciprocal walls were earlier also referred as irreversible ones Janovec81 ; Janovec .
@@
Any nominally charged ferroelectric domain wall is a non-reciprocal wall. In a lamellar structure, the subsequent domains usually have an opposite normal component of the spontaneous polarization and subsequent domain walls then differ by the sign of the interfacial charge. They can be thus denoted as head-to-head (HH) or tail-to-tail (TT) walls, respectively, and the absence of symmetry relation is obvious.
@@
In this paper, we are addressing the question whether and when one can encounter also manifestations of the non-reciprocity of nominally neutral ferroelectric walls. In particular, we explore the case of the so-called Marton2010 R180($\bar{1}\bar{1}2$) walls and argue that these in BiFeO3 experimentally reported experiment but so far little investigated walls are indeed showing remarkable non-reciprocal effects.
@@
BiFeO3 is a well-studied multiferroic material with coexisting ferroelectric, ferroelastic, antiferrodistortive and antiferromagnetic orders and promising applications in spintronics, nonvolatile memory devices, catalysis and sensors wang_2020 ; chaudron_2024 ; YangAbove ; Amdouni ; Huyan2019 . At room temperature, it has rhombohedral $R3c$ symmetry with a spontaneous polarization of about 1 C/m2 and with oxygen octahedra alternatively tilted by about $14^{\circ}~{}$ around the polarization axis lebeugle_2007 ; Catalan09 . The conventional hexagonal unit cell is depicted in Fig. 1a. When neglecting the spontaneous strain, its lattice vectors can be expressed in the basis of the parent high-temperature $Pm\bar{3}m$ phase as ${\bf a_{\rm h}}=(0,1,\bar{1})$, ${\bf b_{\rm h}}=(\bar{1},0,1)$ and ${\bf c_{\rm h}}=(2,2,2)$. This relation allows us to describe crystallographic directions with Miller indices $h,k,l$ of the parent cubic crystal structure. It also implies the existence of the eight equivalent ferroelectric domain states with polarization along $\langle 111\rangle$ cubic diagonals. Corresponding ferroelectric domain walls are primarily distinguished as 180, 109 and 71 ones according to the angles between the polarization in the adjacent domains. Structure and properties of domain walls in BiFeO3 have been subject of multiple experimental Farokhipoor ; Sando14 ; Bhatnagar ; Geng2020 ; Chauleau2020 ; Fusil2022 ; Condurache and theoretical studies Dieguez2013 ; Ren2013 ; ourPRL ; Chauleau2020 ; Lubk ; Wang2013 ; Xue2014 ; Zhang2022 ; Mangeri ; Eliseev . The benchmark systematic first-principles study of Ref. Dieguez2013, explored almost all types of mechanically and electrically compatible (neutral) domain walls. However, non-reciprocal walls were not addressed there.
@@
To demonstrate and explain several remarkable non-reciprocal domain wall effects in ${}^{\circ}~{}$ domain walls of BiFeO3 we consider two antiparallel domain states with polarization along $[111]$ and $[\bar{1}\bar{1}\bar{1}]$. The mechanically and electrically compatible ${}^{\circ}~{}$ domain walls connecting these two states should be parallel to the $[111]$ direction. There are two high-symmetry options complying with this requirement. One option is realized by $(1\bar{1}0)$-type planes, i.e. walls parallel to the mirror symmetry plane of the parent $Pm\bar{3}m$ reference phase. For example, it would be $y^{\prime}z^{\prime}$ plane within the Cartesian coordinate system $x^{\prime}$, $y^{\prime}$ and $z^{\prime}$ with $x^{\prime}\parallel{\bf a_{\rm h}}$ and $z^{\prime}\parallel{\bf c_{\rm h}}$ shown in Fig. 1. The other option represents the ($\bar{1}\bar{1}2$) plane, passing through intersecting 3-fold and 2-fold rotational axes of the parent $Pm\bar{3}m$ structure. Most of the studies have so far considered only the former type. However, as we shall show, the latter ($\bar{1}\bar{1}2$)-type walls, such as DW1 and DW2 depicted in Fig. 1b, are even more important for BiFeO3.
Our predictions are based on an interatomic potential with parameters fitted to ab-initio calculations Graf2014 ; Graf2014b . This interatomic potential describes each atom with a core and a shell with a mutual anharmonic core-shell interaction mimicking the atomic polarizability Tinte2004 . It has been successfully applied to BiFeO3 in few other studies Graf2015 ; Graf2016 ; ourPRL . The domain wall structures were relaxed using classical molecular dynamics simulations under constant stress and temperature conditions, using a supercells with typically 20$\times$20$\times$2 conventional rhombohedral BiFeO3 cells of Fig. 1a (i.e. 24000 atoms) subjected to periodic boundary conditions. A temperature of 1 K was chosen to facilitate effective relaxation while introducing a minimal amount of thermal noise. The simulation began with a 10 ps thermalization, followed by an additional 10 ps during which the average configuration was computed. The molecular dynamics calculations were performed using the DL_POLY software Todorov2006 . Local dipole moments were evaluated for each 5-atom perovskite cell centered in the Bi atoms, based on the positions and charges of cores and shells. For a given cell, one Bi atom with its surrounding twelve O atoms (with a weight of 0.25) and eight coordinating Fe atoms (0.125 weight) were considered. The desired ($\bar{1}2\bar{1}$) domain wall orientation was selected by imposing the suitable step-like distribution of the initial Bi ion off-centering displacements in our simulation box. The periodic boundary conditions requires to introduce two such domain walls, as indicated in Fig. 1b.
@@
After the relaxation, most of the unit cells have their structure similar to that of the homogeneous ground state. For the detailed inspection of the region near the domain wall, polarization of each unit cell was projected on the rotated coordinate system, where $s\parallel[\bar{1}2\bar{1}]$ is perpendicular to the wall, $z^{\prime}\parallel[111]$ is along the polarization, and $t\parallel{\bf b_{\rm h}}$ is perpendicular to the both $s$ and $z^{\prime}$. Resulting polarization profiles in the vicinity of domain wall DW1 and DW2 are shown in Fig. 2a,b. Analogous profiles of the staggered oxygen octahedra tilt rotations $R_{i}$ are shown in Fig. 2c,d. As expected, $P_{\rm s}$ and $P_{\rm t}$ components are zero or close to zero, while the $P_{\rm z^{\prime}}$ component is switching between the nominal spontaneous polarization values. The absence of switching of the $R_{\rm i}$ components confirms that neither DW1 nor DW2 coincide with an anti-phase boundary of the underlying staggered tilt pattern and thus both walls are purely ferroelectric domain walls.
@@
While the profiles of $P_{\rm s}$, $P_{\rm z^{\prime}}$, $R_{\rm s}$, and $R_{\rm z^{\prime}}$ components are same at any line passing through the given wall, the values of $P_{\rm t}$ and $R_{\rm t}$ alternate between two mutually opposite values when comparing lines related by the fractional translation vector ${\bf c_{\rm h}}/2$. It means that the Bloch polarization component averaged over the two lines is zero, but at the same time, there is a staggered (anti-ferroelectric) Bloch component within the wall. In case of the tilt profile, the double signs of $R_{\rm t}$ implies that the Bloch component of the staggered oxygen octahedra rotation averaged over the two octahedra related by ${\bf c_{\rm h}}/2$ is zero, but both octahedra are slightly rotated in the same sense around the $t$-axis.
@@
There are also important differences between DW1 and DW2. First, the DW1 wall thickness defined by slope at the center of the wall as in Refs. Hlinka2006, ; Marton2010, is about 0.7 nm, while the DW2 thickness is only about 0.35 nm. Second, the magnitude of the tilt modulations is much stronger in DW1. Third, the signs of the $R_{\rm s}$ are opposite in DW1 and DW2. A more detailed analysis of these differences within a continuous Ginzburg-Landau theory is a subject of a follow-up paper.
@@
To verify that the profiles of Fig. 2 are the lowest energy domain wall configurations, we have slightly varied the initial positions of the walls, size of the simulation box and we also performed annealing at higher temperatures. In overall, the domain characteristics described above proved to be robust and representative ones.
Nevertheless, an interesting phenomenon was encountered in the course of gradual temperature increase in steps of 20 K up to 300 K extradetails . While the time-averaged polarization map at $z^{\prime}=0$ was practically the same till $T=120$ K (see Fig. 3a), as soon as the temperature reached $T=140$ K, the DW2 started to display significant fluctuations. The original planar wall changed to an irregularly corrugated shape and the $P_{\rm t}$ components started to be disordered. The slow fluctuations are apparent even in the 10 ps averaged polarization maps (see Fig. 3b). Finally, at $T=240$ K, the shape of DW2 abruptly transformed into a zigzag wall with ${}^{\circ}~{}$ folds. This form of DW2 then persists till $T=300$ K (see Fig. 3c) and it remains there even when the system is cooled back down to $T=1$ K.
@@
Finally, we also tried to stabilize small diameter columnar nanodomain of inverted polarization. The only encountered stable nanodomain had a shape of a right prism with a equilateral triangular cross section as shown in Fig. 4. We could see that one of the sides of this triangular prism is always parallel the walls of Fig. 1b, and that the wall profiles correspond to that of DW1 in Fig. 2a,c. We have never observed a spontaneous formation of DW2 walls or $(1\bar{1}0)$-type walls, suggesting that DW1 is actually the lowest energy variant of the 180 domain wall.
@@
Summarizing, we have shown that the DW2 differs from DW1 by the thickness, by the amplitudes of the in-wall order parameters and also by the stability. The ensemble of all these findings implies that DW1 and DW2 are not related by any symmetry operation and so these walls are clearly non-reciprocal ones.
@@
Figure 4: Stable nanodomains of antiparallel polarization. (a) Polarization map at $z^{\prime}=0$ cross section through a stable prismatic nanodomain. (b) The excess energy due to the domain walls as a function of wall-fold distance $l$ and the linear fit to the data.
@@
To facilitate the discussion, let us introduce a more precise domain wall symbol $[A|{\bf{n}}\rangle B]$. Here $A$ and $B$ stand for domain states, while the oriented domain wall normal $\bf{n}$ points in the real space towards the domain that is ascribed to the state of the right-hand side of the symbol – as it is mnemonically indicated by the embedded “ket” symbol. With the already introduced cubic Miller indices for spontaneous polarization directions and oriented domain wall normals, our DW1 of Fig. 1b can be described as $[111|1\bar{2}1\rangle\bar{1}\bar{1}\bar{1}]=[\bar{1}\bar{1}\bar{1}|\bar{1}2\bar{1}\rangle 111]$, while DW2 can be described as $[111|\bar{1}2\bar{1}\rangle\bar{1}\bar{1}\bar{1}]=[\bar{1}\bar{1}\bar{1}|1\bar{2}1\rangle 111]$.
@@
Symmetry of planar domain walls was systematically developed in Refs. Janovec81, ; Janovec, . The basic assumption is that specific symmetries are described by certain subgroups of the parent phase crystallographic group $G$. The symmetry of the $[A|{hkl}\rangle B]$ wall in the usual sense is described by a group $T_{AB}|{hkl}\rangle$, composed of those parent phase symmetry group operations, that generates image structures indistinguishable from the original domain wall structure. It is a crystallographic group with two-dimensional translational symmetry. The highest possible symmetry of the wall $T_{AB}^{0}(hkl)$ can be determined considering the orientation of the domain wall plane and its centering in the crystal structure Janovec81 ; Janovec ; Schranz . In the present case, the $2_{\rm t}/m_{\rm t}$ factor group of $T_{AB}^{0}(hkl)$ contains the identity, inversion $i$, mirror $m_{\rm t}$ perpendicular to $t$ axis, and the 2-fold rotation $2_{\rm t}$ (parallel to the $t$ axis). The profiles of Fig. 2 agree with this symmetry within the precision of our numerical calculations.
@@
The non-reciprocity phenomenon is related to the symmetry group of domain wall pair $\bar{J}_{AB}(hkl)$, which collects all parent phase symmetry operations that preserve or exchange the domain states and preserve the domain wall plane $(hkl)$. It is a supergroup of $T_{AB}^{0}(hkl)$ because it may also contain operations that interchange the sides of the wall but do not alter the domain state (pure side-reversal operations) or the operations that reverse the domain states but do not interchange sides of the wall (pure state-reversal operations). The existence of such pure side-reversal or pure state-reversal symmetry operations would imply the reciprocity (reversibility) of the wall. In the case of the present wall, no such symmetry operations exists and $\bar{J}_{AB}(hkl)$ equals to $T_{AB}^{0}(hkl)$.
@@
Finally, the group of the unordered domain pair, $J_{AB}$, involves all operations of $G$ maintaining or interchanging the two domain states Schranz . In our case the $\bar{3}m$ factor group of $J_{AB}$ is a nontrivial supergroup of the factor group of $\bar{J}_{AB}|(hkl)$ what implies that there are other walls equivalent to DW1 that join the same two domain states. It can be seen that these are the $[111|\bar{2}11\rangle\bar{1}\bar{1}\bar{1}]=[\bar{1}\bar{1}\bar{1}|2\bar{1}\bar{1}\rangle 111]$ and the $[111|11\bar{2}\rangle\bar{1}\bar{1}\bar{1}]=[\bar{1}\bar{1}\bar{1}|\bar{1}\bar{1}2\rangle 111]$ walls corresponding to the segments of the the zigzag wall in Fig. 3b. All these three equivalent walls are realized in the three sides of the triangular prism in Fig. 4a.
@@
The non-reciprocal walls should also differ in their planar energy density. The irreversible decomposition of DW2 wall documented in Fig. 3 and the systematic preference for DW1-type domain walls in stable prismatic nanodomains (Fig. 4a) suggests that the planar domain wall energy density $\sigma_{2}$ of DW2 is significantly higher then $\sigma_{1}$ of DW1. The broadly applied procedure of the domain wall energy density determination from computational experiments consists of (i) calculation of energy of planar domains with desired crystallographic orientation in a periodic cell like that of Fig. 1, (ii) subtracting the corresponding energy of the single domain state, and (iii) dividing the difference by the total surface of the walls within the supercell. In case of non-reciprocal wall pair, however, this yields only the average energy density $(\sigma_{1}+\sigma_{2})/2$. In order to overcome this topological obstacle, we have thus used the prismatic nanodomain of Fig. 4a. As argued above, all three sides walls are equivalent to DW1 wall. For elimination of the domain fold energy contributions, we have calculated a whole sequence of relaxed nanodomain configurations with the wall-fold distance (the arm of the displayed triangle) $l$, ranging from 5 to 20 nm. The slope of the excess energy associated with the domain walls vs the distance $l$ shown in Fig. 4b yielded $\sigma_{1}\approx 40$ mJ/m2. In combination with the average energy density obtained from the standard infinite wall configuration of Fig. 1, we could see that $\sigma_{2}\approx 130$ mJ/m2, which is indeed significantly larger than $\sigma_{1}$.
Why despite many prior computational studies of domain walls the phenomenon of non-reciprocity escaped from the attention so far? One of the reasons is that the only non-reciprocal and highly symmetric wall in perovskite ferroelectrics is this ($\bar{1}\bar{1}2$)-oriented 180 wall of rhombohedral phase. Moreover, embedding of ($\bar{1}\bar{1}2$) oriented wall requires a somewhat larger periodic supercell. The second reason is that the lowest-order coupling term introducing the non-reciprocity in Landau-Ginzburg theory is the third power of polarization gradient. This term has been systematically omitted in the earlier works. At the same time, without a specific structural reason, the non-reciprocal effects can be quite small. In case of BiFeO3, the extraordinary enhancement mechanism consists in the coupling of the polarization to the oxygen octahedra tilt distortions by a third-order mixed polarization-tilt gradient term, combined with a Lifshitz-type bilinear coupling term tobe .
@@
In summary, we have explored the 180 domain walls of BiFeO3 and find out that so far little investigated ($\bar{1}\bar{1}2$)-oriented walls are the lowest energy among the R180 ones. We have given computational and symmetry arguments why these ($\bar{1}\bar{1}2$)-oriented walls are non-reciprocal and explained various profound consequences of this peculiar property.
@@
This work was supported by the Czech Science Foundation (project no. 19-28594X) and by the European Union’s Horizon 2020 research and innovation programme under grant agreement no. 964931 (TSAR). MAPG acknowledge the European Union and the Czech Ministry of Education, Youth and Sports (Project: MSCA Fellowship CZ FZU I - CZ.02.01.01/00/22__\__010/0002906). Computational resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic.